OpenStudy (anonymous):

A constant force F = −4i + 1j − 10k moves an object along a straight line from point (4, -3, 7) to point (-10, 9, 7). Find the work done if the distance is measured in meters and the magnitude of the force is measured in newtons. Work: ??? Nm

OpenStudy (anonymous):

you have to parameterize the line first the do the line integral from [0,1] if you parameterize the line right then it will be [0,1] *r'(t)dt onece you parametrize you take derivitve d/dt and take the magnitude. that will turn the ds=dt the dot F with r'(t) and your ||r'(t)|| will be a constant so pull that to the front , the integrate.

OpenStudy (anonymous):

wow ! that's a hard one to do ! do u mind showing more detail please ! perhaps numerical please

OpenStudy (anonymous):

W=F*distance or ds which from calc 1 or 2 =sqrt(i component squared+....+z component squared)

OpenStudy (anonymous):

i am going to the gym right now, when i come back if someone has not yet answered it i will do it. go to this website in the mean time.... http://tutorial.math.lamar.edu/Classes/CalcIII/LineIntegralsPtI.aspx

OpenStudy (anonymous):

so sqrt of 117 ! times what then !?

OpenStudy (anonymous):

http://www.ltcconline.net/greenl/courses/202/vectorIntegration/lineIntegrals.htm then go half way done they do an excellent job of explaining Work=F*r'(t)dt

OpenStudy (anonymous):

but i got lost half way it explains very good though ! too much computation

OpenStudy (anonymous):

(-10-4)it+(9--3)jt+(7-7)kt x=4-14t y=-3+12t z=7 tE[0,1] r(t)=-14t,12t,0 r'(t)=-14,12,0 ||r'(t)||=sqrt(196+144)=sqrt(340) Line Integral [0,1]=(-4,1,-10)*(-14,12,0)sqrt(340)dt Line integral [0,1]56+12=68(sqrt(340))dt which to me looks like 68sqrt(340) 340=34*10=17*2*2*5 2sqrt(85)*68=136sqrt(85) Tell me if thats what you see in your book

OpenStudy (anonymous):

Do you know how to parameterize a line?

OpenStudy (anonymous):

no !

OpenStudy (anonymous):

if p=(a,b,c) and Q=(x,y,z) PQ=(x-a)it+(y-b)jt+(z-c)kt this is also your r(t) equation =P+PQ so x=a+(x-a)t y=b+(y-b)t z=c+(z-c)t

OpenStudy (anonymous):

i tried 136sqrt(85) its not correct !

OpenStudy (anonymous):

that's why i hate online homework ! grrrr....

OpenStudy (anonymous):

ok lemme go to the gym, if anyone else answer im sorry i just did this all on the comp no scratch work. Ill be back less than 2 hrs ill search for you

OpenStudy (anonymous):

oh online HW ios horrible im sorry

OpenStudy (anonymous):

have fun man and thanks again

OpenStudy (anonymous):

your welcome.

OpenStudy (anonymous):

Try your answer with out the SQRT(340)

OpenStudy (anonymous):

nope !

OpenStudy (anonymous):

ok ill do it i swear add me on Facebook too ill send you a step by step thing of how to do it. names brett cory and facebook pic is orange

OpenStudy (anonymous):

its due in 2 hours though !

OpenStudy (anonymous):

new port beach !?

OpenStudy (anonymous):

yeah. F=-4i+1j-10k r'(t)=-14i,12j,0k F*r'(t)=56+12+0=68dt INT[0,1]68dt=68mN stil;l no?

OpenStudy (anonymous):

it is 68 :D thanks man

OpenStudy (anonymous):

just added u on fb too name is Ehsan Rahimzadegan

OpenStudy (anonymous):

AWWW ok glad to help you can still ad me n ask calc 3 questions if you would like

OpenStudy (anonymous):

yeah man thanks i will !